# Arbitrage-free catastrophe reinsurance valuation for compound dynamic contagion claims

### v1

@@ 2.3 Time-inhomogeneous compound dynamic contagion process
For the time-inhomogeneous CDCP, the infinitesimal generator ${\cal A}$ of the joint process $\{\left(\lambda_{t},N_{t},C_{t},{M_{t}},{\Lambda_{t}},t\right)\}_{t\in{\cal T}}$ acting on a function $f\left(\lambda,n,c,{m},{\Lambda},t\right)\in\mathcal{D}\left(\mathcal{A}\right)$ is given by:
$\mathcal{A}f\left(\lambda,\Lambda,n,c,{m},{\Lambda},t\right)$
@@
$\quad+\lambda\left[\int_{0}^{\infty}\int_{0}^{\infty}f\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}G(y;t)\mathop{}\!\mathrm{d}J(\xi)-f\left(\lambda,n,c,{m},{\Lambda},t\right)\right]$
$\quad+\rho(t)\left[\int_{0}^{\infty}f\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}H(x;t)-f\left(\lambda,n,c,{m},{\Lambda},t\right)\right],$ (2.7)
where $\mathcal{D}\left(\mathcal{A}\right)$ is the domain of the generator $\mathcal{A}$ such that $f\left(\lambda,n,c,{m},{\Lambda},t\right)$ is differentiable with respect to $\lambda$ and $t$, and
$\left|\int_{0}^{\infty}\int_{0}^{\infty}f\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}G(y;t)\mathop{}\!\mathrm{d}J(\xi)-f\left(\lambda,n,c,{m},{\Lambda},t\right)\right|<\infty\text{,}$
$\left|\int_{0}^{\infty}f\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}H(x;t)-f\left(\lambda,n,c,{m},{\Lambda},t\right)\right|<\infty\text{.}$
Setting $\mathcal{A}$ $f\left(\lambda,\Lambda,n,c,t\right)=\lambda$ in Section 2.3, we have
@@
As $\lambda_{t}-\lambda_{0}-\int_{0}^{t}\mathcal{A}\lambda_{s}\mathop{}\!\mathrm{d}s$ is an $\mathcal{F}_{t}$-martingale, we have
$\mathbb{E}\Biggl{(}\lambda_{t}-\int_{0}^{t}\mathcal{A}\lambda_{s}\mathop{}\!\mathrm{d}s\mid\lambda_{0}\Biggr{)}=\lambda_{0}.$
Hence,
@@
and solving this ODE, we have solution of Eq. 2.4. Using Eq. 2.4, given $N_{0}=C_{0}=0$, we can derive
@@
The expressions for Eqs. 2.5 and 2.6 would be very long with various simple exponential functions. To save space, we just leave their concise expressions, but we will compare them with corresponding time-homogeneous cases in the context of arbitrage-free catastrophe reinsurance premiums in Section 4. Appendix A contains the algorithm we developed to simulate from this new process.
@@ 3 Insurance market and no-arbitrage
We consider a liquid reinsurance market in which an insurer can cede part of its aggregate risk to a reinsurer via a reinsurance arrangement at any time. Here the no-arbitrage valuation approach is adopted to value a reinsurance contract. Sondermann (1991) pioneered the no-arbitrage valuation approach to reinsurance contracts. Dassios and Jang (2003) proposed the use of the Esscher transform to specify an equivalent martingale measure for valuing reinsurance contracts by the no-arbitrage valuation principle. Jang and Krvavych (2004) introduced the use of the Esscher transform to determine arbitrage-free premiums for extreme losses. The developments here follow those in, for example, Sondermann (1991), Dassios and Jang (2003) and Jang and Krvavych (2004).
@@
For each $t\in{\cal T}$, a reinsurance strategy starting at time $t$ is an $\mathbb{F}$-predictable process $\{\phi_{u}|u\in[t,T]\}$ on $(\Omega,{\cal F},\mathbb{P})$ such that for all $u\in[t,T]$, $\phi_{u}\in[0,1]$. Write ${\cal H}_{t}$ for the space of all reinsurance strategies starting at time $t$.
@@
For a given time $t\in{\cal T}$, if the insurer selects a reinsurance strategy $\{\phi_{u}|u\in[t,T]\}\in{\cal H}_{t}$ at time $t$, then the insurer’s final gain at time $T$ is given by:
@@
Following Sondermann (1991), a definition of an arbitrage (reinsurance) strategy is provided.
@@
For each $t\in{\cal T}$, a reinsurance strategy starting at time $t$, say $\{\phi_{u}|u\in[t,T]\}\in{\cal H}_{t}$, is said to be an arbitrage reinsurance strategy if the respective insurer’s final gain at time $T$ satisfies
@@
Intuitively, an arbitrage reinsurance strategy is a reinsurance strategy that allows for a possible profit without the possibility of a loss.
As noted in, for example, Dassios and Jang (2003), using the fundamental theorem of asset pricing (Harrison and Kreps (1979) and Harrison and Pliska (1981)), the insurance business characterized by the surplus process $\{R_{t}\}_{t\geq{\cal T}}$ does not allow arbitrage opportunities if there is “essentially” an equivalent martingale measure $\mathbb{P}^{\ast}$ such that the process $\{R_{t}\}_{t\in{\cal T}}$ is an $(\mathbb{F},\mathbb{P}^{*})$-martingale.
In fact, the insurance market is incomplete. Consequently, there is more than one equivalent martingale measure. As in, for example, Dassios and Jang (2003) and Jang and Krvavych (2004), the Esscher transform will be employed in Section 4 to change probability measures. Consequently, an equivalent martingale measure can be selected. The Esscher transform provides a convenient way to determine an equivalent martingale measure. The seminal work by Gerber and Shiu (1994) introduced the use of the Esscher transform to option pricing. Bühlmann et al. (1996) and Kallsen and Shiryaev (2002) generalized the use of the Esscher transform to select an equivalent martingale measure in a general semimartingale market. Since then, the use of the Esscher transform in option valuation has been widely studied. See, for example, Gerber and Shiu (1994), Gerber and Shiu (1996), Dassios and Jang (2003), Jang and Krvavych (2004), Siu et al. (2004), Elliott et al. (2005), Siu (2005), Goovaerts and Laeven (2008), Badescu et al. (2009), Eberlein et al. (2009), amongst many others. In the following, we recall the definition from Dassios and Jang (2003), where the Esscher transform was used to select an equivalent martingale measure relevant to arbitrage-free reinsurance strategies in an (incomplete) insurance market. The definition is standard and is presented here for the sake of completeness.
@@ 3.1 Equivalent martingale measures
In this section, we extend the change of measure for the Cox process with shot-noise intensity (Dassios and Jang, 2003) to that for a standard dynamic contagion process (Dassios and Zhao, 2011).
@@
A probability measure $\mathbb{P}^{\ast}$ is said to be an equivalent martingale probability measure (i.e., $\mathbb{P}^{\ast}$ is equivalent to $\mathbb{P}$ on ${\cal F}_{T}$) if it satisfies the following three properties:
@@
    The Radon–Nikodym derivative $\frac{d\mathbb{P}^{\ast}}{d\mathbb{P}}\in L^{2}(\Omega,{\cal F}_{T},\mathbb{P})$, where $L^{2}(\Omega,{\cal F}_{T},\mathbb{P})$ is the space of square-integrable, ${\cal F}_{T}$-measurable, random variables on $(\Omega,\mathbb{P})$.
@@
    The surplus process $\{R_{t}\}_{t\in{\cal T}}$ is an $(\mathbb{F},\mathbb{P}^{\ast})$-martingale. In particular,
@@
    for any $0\leq s\leq t\leq T$, where $\mathbb{E}^{\ast}[\cdot|{\cal F}_{s}]$ denotes the conditional expectation given ${\cal F}_{s}$ with respect to $\mathbb{P}^{\ast}$.
Definition 3.4 (Esscher transform).
Let $\{X_{t}\}_{t\in{\cal T}}$ be a stochastic process on the complete filtered probability space $(\Omega,{\cal F},\mathbb{F},\mathbb{P})$ such that the exponential process $\{\mathrm{e}^{\hbar X_{t}}\}_{t\in{\cal T}}$ is an $(\mathbb{F},\mathbb{P})$-martingale, for some $\hbar\in\mathbb{R}$. For a measurable and integrable function $f:\mathbb{R}\to\mathbb{R}$, the expectation of $f(X_{t})$ with respect to an Esscher probability measure $\mathbb{P}^{\ast}$ corresponding to the Esscher transform with the Esscher parameter $\hbar$ is given by:
$\mathbb{E}^{\ast}[f(X_{t})]=\mathbb{E}\left[f(X_{t})\times\frac{\mathrm{e}^{\hbar X_{t}}}{\mathbb{E}\left[\mathrm{e}^{\hbar X_{t}}\right]}\right]=\frac{\mathbb{E}\left[f(X_{t})\times\mathrm{e}^{\hbar X_{t}}\right]}{\mathbb{E}\left[\mathrm{e}^{\hbar X_{t}}\right]},$ (3.2)
provided that $\mathbb{E}\left[\mathrm{e}^{\hbar X_{t}}\right]<\infty$.
Indeed, Eq. 3.2 follows by a version of the Bayes’ rule and setting the Radon–Nikodym derivative for the measures change as the Esscher transform with the Esscher parameter $\hbar$. From Definition 3.4, to determine an equivalent martingale probability measure $\mathbb{P}^{\ast}\sim\mathbb{P}$ on ${\cal F}_{T}$, we need to find an $(\mathbb{F},\mathbb{P})$- exponential martingale $\{\mathrm{e}^{\hbar X_{t}}\}_{t\in{\cal T}}$ such that it is used to define the Radon–Nikodym derivative $\frac{d\mathbb{P}^{\ast}}{d\mathbb{P}}$ for changing the probability measures from the original probability measure $\mathbb{P}$ to an equivalent martingale probability measure $\mathbb{P}^{\ast}$. That is, we need to find $\hbar$ so that $\{\mathrm{e}^{\hbar X_{t}}\}_{t\in{\cal T}}$ is an $(\mathbb{F},\mathbb{P})$-exponential martingale. Then, we set $\frac{d\mathbb{P}^{\ast}}{d\mathbb{P}}=\frac{\mathrm{e}^{\hbar X_{t}}}{\mathbb{E}\left[\mathrm{e}^{\hbar X_{t}}\right]}$ (i.e., the Esscher transform with the Esscher parameter $\hbar$). To do this, we start with Theorem 3.1 giving a martingale which will be used to specify the Radon–Nikodym derivative for the measures change via the Esscher transform.
@@
Suppose that $\rho>0$ for all $t\in{\cal T}$. For constants $\theta,{\psi},\nu$, we have an $(\mathbb{F},\mathbb{P})$-martingale
@@
where $^{\hat{h}}(\cdot)$, ${\hat{g}}(\cdot)$ and ${\hat{j}}(\cdot)$ are the Laplace transforms defined by Eq. 2.3.
@@
Firstly, we consider a function $f\left(\lambda_{t},N_{t},C_{t},{M_{t}},{\Lambda_{t}},t\right)$ defined by:
$f\left(\lambda_{t},N_{t},C_{t},{M_{t}},{\Lambda_{t}},t\right)~{}=~{}\mathrm{e}^{K(t)}\theta^{N_{t}}\mathrm{e}^{B(t)\lambda_{t}}~{}\mathrm{e}^{-\nu C_{t}}{\mathrm{e}^{\phi\Lambda_{t}}\psi^{M_{t}}},$
@@
[…]
$f\left(\lambda,\Lambda,n,c,{m},{\Lambda},t\right)=\mathrm{e}^{K(t)}\theta^{n}\mathrm{e}^{B(t)\lambda}~{}\mathrm{e}^{-\nu c}{\mathrm{e}^{\phi\Lambda}\psi^{m}},$
@@
00 0 $=K^{\prime}(t)+B^{\prime}(t)\lambda+\delta\left(a-\lambda\right)B(t)+\lambda\phi$
@@
00 0 $=\lambda\left\{B^{\prime}(t)-\delta B(t)+\phi+\theta\int_{0}^{\infty}\int_{0}^{\infty}\mathrm{e}^{B(t)y}\mathrm{e}^{-\nu\xi}\mathop{}\!\mathrm{d}G(y)\mathop{}\!\mathrm{d}J(\xi)-1\right\}$
@@
00 0 $=\lambda\bigg{\{}B^{\prime}(t)-\delta B(t)+\phi+\theta\hat{j}(\nu)\hat{g}(-B(t))-1\bigg{\}}+K^{\prime}(t)+\delta aB(t)+\rho\left[\psi\hat{h}(-B(t))-1\right],$
@@
    $\mathcal{G}(B):=\int^{B}_{b}\frac{du}{f_{1}(u)},\qquad B\in\left(b,B^{+}\right),$
@@
$\frac{\mathop{}\!\mathrm{d}f_{1}(B)}{\mathop{}\!\mathrm{d}B}\mid_{B=0}=\delta-\theta\hat{j}(\nu)\mu_{G}>0.$ (3.6)
@@
$\frac{\mathop{}\!\mathrm{d}^{2}f_{1}(B)}{\mathop{}\!\mathrm{d}B^{2}}=-\theta\hat{j}(\nu)\int^{\infty}_{0}\mathrm{e}^{uy}y^{2}\mathop{}\!\mathrm{d}G(u)<0.$
Then $f_{1}(B)$ is strictly concave. This, together with Eqs. 3.5 and 3.6, imply that $f_{1}(B)>0$, for $B\in\left(0,B^{+}\right)$, where $B^{+}>0$ is the smallest positive solution to $f_{1}(B)=0$.
Taking $B=u$ and $t=\tau$, Eqs. 3.3 and 3.4 lead to:

### v2

@@ 2.2.2 Time-inhomogeneous compound dynamic contagion process
The infinitesimal generator $\cal A$} of the joint process $\{\left(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t\right)\}_{t\in{\cal T}}$ acting on a function $f(\lambda,n,c,m,\Lambda,t)\in\mathcal{D}\left(\mathcal{A}\right)$ for the time-inhomogeneous CDCP is given by:
$\hskip-14.22636pt\mathcal{A}f(\lambda,n,c,m,\Lambda,t)$
@@
$\quad+\lambda\left[\int_{0}^{\infty}\int_{0}^{\infty}f(\lambda+y,n+1,c+\xi,m,\Lambda,t)\mathop{}\!\mathrm{d}G(y;t)\mathop{}\!\mathrm{d}J(\xi)-f(\lambda,n,c,m,\Lambda,t)\right]$
$\quad+\rho(t)\left[\int_{0}^{\infty}f(\lambda+x,n,c,m+1,\Lambda,t)\mathop{}\!\mathrm{d}H(x;t)-f(\lambda,n,c,m,\Lambda,t)\right],$ (2.9)
where $\mathcal{D}\left(\mathcal{A}\right)$ is the domain of the generator $\mathcal{A}$ such that $f(\lambda,n,c,m,\Lambda,t)$ is differentiable with respect to $\lambda$ and $t$, and
$\left|\int_{0}^{\infty}\int_{0}^{\infty}f(\lambda+y,n+1,c+\xi,m,\Lambda,t)\mathop{}\!\mathrm{d}G(y;t)\mathop{}\!\mathrm{d}J(\xi)-f(\lambda,n,c,m,\Lambda,t)\right|<\infty\text{,}$
$\left|\int_{0}^{\infty}f(\lambda+x,n,c,m+1,\Lambda,t)\mathop{}\!\mathrm{d}H(x;t)-f(\lambda,n,c,m,\Lambda,t)\right|<\infty\text{.}$
Setting $f(\lambda,n,c,m,\Lambda,t)=\lambda$ in Section 2.2.2, we have
@@
For each $t\geq 0$, ${\lambda}^{\dagger}_{t}:=\lambda_{t}-\lambda_{0}-\int_{0}^{t}\mathcal{A}\lambda_{s}\mathop{}\!\mathrm{d}s$. Then for any $s<t$, we have:
$\mathbb{E}[\lambda^{\dagger}_{t}-\lambda^{\dagger}_{s}\,\mid\,{\cal F}_{s}]=\mathbb{E}\left[\lambda_{t}-\lambda_{s}-\int_{s}^{t}\mathcal{A}\lambda_{u}\mathop{}\!\mathrm{d}u\,\mid\,{\cal F}_{s}\right],\quad\mbox{$\mathbb{P}$-a.s.}$
By the Dynkin formula, since $t<\infty$,
$\mathbb{E}[\lambda_{t}\,|\,{\cal F}_{s}]=\lambda_{s}+\mathbb{E}\left[\int_{s}^{t}\mathcal{A}\lambda_{u}\mathop{}\!\mathrm{d}u\ |\ {\cal F}_{s}\right],\quad\mbox{$\mathbb{P}$-a.s.}$
Consequently, by Fubini’s theorem,
$\mathbb{E}[\lambda^{\dagger}_{t}-\lambda^{\dagger}_{s}\,|\,{\cal F}_{s}]=\mathbb{E}\left[\lambda_{t}-\lambda_{s}-\int_{s}^{t}\mathcal{A}\lambda_{u}\mathop{}\!\mathrm{d}u\,\mid\,{\cal F}_{s}\right]=0,\quad\mbox{$\mathbb{P}$-a.s.}$
This implies that
$\mathbb{E}[\lambda^{\dagger}_{t}\,|\,{\cal F}_{s}]=\lambda^{\dagger}_{s},\quad\mbox{$\mathbb{P}$-a.s.}$
Since $\{\lambda^{\dagger}_{t}\}_{t\geq 0}$ is an $(\mathbb{F},\mathbb{P})$-martingale, we must have:
$\mathbb{E}[\lambda^{\dagger}_{t}\,|\,{\cal F}_{0}]=\mathbb{E}[\lambda^{\dagger}_{t}\,|\,\lambda_{0}]=\mathbb{E}\Biggl(\lambda_{t}-\lambda_{0}-\int_{0}^{t}\mathcal{A}\lambda_{s}\mathop{}\!\mathrm{d}s\,\mid\,\lambda_{0}\Biggr)=0,\quad\mbox{$\mathbb{P}$-a.s.}$
Consequently,
$\mathbb{E}[\lambda_{t}\,|\,\lambda_{0}]=\lambda_{0}+\int^{t}_{0}\mathbb{E}\left[{\cal A}\lambda_{u}\,|\,\lambda_{0}\right]\mathop{}\!\mathrm{d}u,\quad\mbox{$\mathbb{P}$-a.s.}$
Then,
@@
and solving this ODE, we obtain Eq. 2.6 as its solution. Using Eq. 2.6, given $N_{0}=C_{0}=0$, we can derive
@@
The expressions for Eqs. 2.7 and 2.8 would be very long with various simple exponential functions. To save space, we just leave their concise expressions, but we will compare them with corresponding time-homogeneous cases in the context of arbitrage-free catastrophe reinsurance premiums in Section 4. Appendix B contains the algorithm we developed to simulate from this new process.
@@ 3 Insurance market and no-arbitrage
We consider a liquid reinsurance market in which an insurer can cede part of its aggregate risk to a reinsurer via a reinsurance arrangement at any time. Here the no-arbitrage valuation approach is adopted to value a reinsurance contract. Sondermann (1991) pioneered the no-arbitrage valuation approach to reinsurance contracts. Dassios and Jang (2003) proposed the use of the Esscher transform to specify an equivalent martingale measure for valuing reinsurance contracts by the no-arbitrage valuation principle. Jang and Krvavych (2004) introduced the use of the Esscher transform to determine arbitrage-free premiums for extreme losses. It may also be noted that Wüthrich et al. (2010) discussed linear pricing functionals and market consistent valuation for insurance products under the no-arbitrage principle. The developments here follow those in, for example, Sondermann (1991), Dassios and Jang (2003) and Jang and Krvavych (2004).
@@
For each $t\in{\cal T}$, a reinsurance strategy starting at time $t$ is an $\mathbb{F}$-predictable process $\{\phi_{u}\mid u\in[t,T]\}$ on $(\Omega,{\cal F},\mathbb{P})$ such that for all $u\in[t,T]$, $\phi_{u}\in[0,1]$. Here, $\phi_{u}$ represents the proportion of the insurer’s liability transferred to the reinsurer at time $u$. Note that $\phi_{u}=0$” corresponds to no reinsurance and that $\phi_{u}=1$” corresponds to full reinsurance. Write ${\cal H}_{t}$ for the space of all reinsurance strategies starting at time $t$.
@@
The assumption of a constant interest rate simplifies the computation of the premium in Section 4.1.
For a given time $t\in{\cal T}$, if the insurer selects a reinsurance strategy $\{\phi_{u}\mid u\in[t,T]\}\in{\cal H}_{t}$ at time $t$, then the insurer’s final gain at time $T$ is given by:
@@
A strategy $\{\phi_{u}\mid u\in[t,T]\}$ that allows for profit with no possibility of loss is called an arbitrage strategy. Hence, following Sondermann (1991), we give the following definition of an arbitrage (reinsurance) strategy.
@@
For each $t\in{\cal T}$, a reinsurance strategy starting at time $t$, say $\{\phi_{u}\mid u\in[t,T]\}\in{\cal H}_{t}$, is said to be an arbitrage reinsurance strategy if the respective insurer’s final gain at time $T$ satisfies
@@
[…]
In fact, the insurance market is incomplete. Consequently, there is more than one equivalent martingale measure. As in, for example, Dassios and Jang (2003) and Jang and Krvavych (2004), the Esscher transform will be employed in Section 4 to change probability measures. Consequently, an equivalent martingale measure can be selected. The Esscher transform provides a convenient way to determine an equivalent martingale measure. The seminal work by Gerber and Shiu (1994) introduced the use of the Esscher transform to option pricing. Bühlmann et al. (1996) and Kallsen and Shiryaev (2002) generalised the use of the Esscher transform to select an equivalent martingale measure in a general semimartingale market. Since then, the use of the Esscher transform in option valuation has been widely studied. See, for example, Gerber and Shiu (1994), Gerber and Shiu (1996), Dassios and Jang (2003), Jang and Krvavych (2004), Siu et al. (2004), Elliott et al. (2005), Siu (2005), Goovaerts and Laeven (2008), Badescu et al. (2009), Eberlein et al. (2009), amongst many others. In the following, we recall the definition from Dassios and Jang (2003), where the Esscher transform was used to select an equivalent martingale measure relevant to arbitrage-free reinsurance strategies in an (incomplete) insurance market. The definition is standard and is presented here for the sake of completeness.
@@ 3.1 Equivalent martingale measures
In this section, we extend the measure change for the Cox process with shot-noise intensity (Dassios and Jang, 2003) to that for a standard dynamic contagion process (Dassios and Zhao, 2011). For further details, we refer to Sondermann (1991), where the no-arbitrage approach to pricing reinsurance contracts was first introduced.
Before introducing the formal notion of an equivalent martingale measure, we emphasize that the arbitrage-free pricing method considered in this paper is understood in the actuarial sense, rather than through replication in a liquid financial market. Specifically, we rule out reinsurance strategies that generate non-negative terminal surplus almost surely with strictly positive expected gain, in accordance with Definition 3.2, and adopt the Esscher transform as a risk-adjusted pricing principle reflecting risk aversion and solvency considerations. In our modeling setup, the market is incomplete and contingent claims such as catastrophe reinsurance contracts cannot be replicated via dynamic hedging; valuation and hedging therefore “divorce”, as noted in Madan and Schoutens (2022) (Chapter 13, Page 171 therein). This makes financial arbitrage arguments based on dynamic replication, such as those in Artzner et al. (2024), less relevant in our setting. Our approach is particularly appropriate for catastrophe reinsurance contracts, which are written over finite (typically annual or multi-year) horizons and are not dynamically hedgeable, and operates within an actuarial martingale valuation paradigm under which the insurer’s surplus process is a martingale under the Esscher-transformed measure (Definition 3.3), thereby ensuring actuarial no-arbitrage pricing for non-traded catastrophic risks.
@@
A probability measure $\tilde{\mathbb{P}}$ is said to be an equivalent martingale probability measure (i.e., $\tilde{\mathbb{P}}$ is equivalent to $\mathbb{P}$ on ${\cal F}_{T}$) if it satisfies the following three properties:
@@
    The Radon–Nikodym derivative $\frac{\mathop{}\!\mathrm{d}\tilde{\mathbb{P}}}{\mathop{}\!\mathrm{d}\mathbb{P}}\in L^{2}(\Omega,{\cal F}_{T},\mathbb{P})$, where $L^{2}(\Omega,{\cal F}_{T},\mathbb{P})$ is the space of square-integrable, ${\cal F}_{T}$-measurable, random variables on $(\Omega,\mathbb{P})$.
@@
    $\tilde{\mathbb{E}}\left[R_{t}\mid\mathcal{F}_{s}\right]=R_{s},\text{ \ }\tilde{\mathbb{P}}-\textit{a.s}.$
    for any $0\leq s\leq t\leq T$, where $\tilde{\mathbb{E}}[\cdot\,|\,{\cal F}_{s}]$ denotes the conditional expectation given ${\cal F}_{s}$ with respect to $\tilde{\mathbb{P}}$.
To simplify the discussion, given that we assume a constant interest rate, we take the interest rate to be zero. The framework extends straightforwardly to a non-zero constant interest rate by working with discounted surplus processes, as detailed in Appendix A. However, the incorporation of a stochastic interest rate, while it is more realistic, would complicate the modeling framework. Consequently, to simplify the discussion, we do not consider the situation of a stochastic interest rate here.
Remark 3.1.
It may be noted that Definition 3.3 (iii) can be expressed as follows: assuming $R_{0}=0$, (i.e., the surplus at the initial time 0 is zero), we have:
$\tilde{\mathbb{E}}\!\left(R_{t}\mid\mathcal{F}_{0}\right)=\tilde{\mathbb{E}}\!\left(P_{t}-C_{t}\mid\mathcal{F}_{0}\right)=\tilde{\mathbb{E}}\!\left(P_{t}\mid\mathcal{F}_{0}\right)-\tilde{\mathbb{E}}\!\left(C_{t}\mid\mathcal{F}_{0}\right)=0.$
Consequently,
$\tilde{\mathbb{E}}\!\left(C_{t}\mid\mathcal{F}_{0}\right)=\tilde{\mathbb{E}}\!\left(P_{t}\mid\mathcal{F}_{0}\right)=\text{an arbitrage-free insurance premium.}$
In the sequel, we shall adopt the Esscher transform to select an equivalent martingale measure. To do this, we start with Theorem 3.1 giving a (local)-martingale which will be considered when specifying the Radon–Nikodym derivative for changing probability measures via the Esscher transform.
@@
Suppose that $\rho>0$ for all $t\in{\cal T}$. For constants $\theta,\nu,\phi,\psi$, we have an $(\mathbb{F},\mathbb{P})$-(local)-martingale
@@
[…]
@@
Firstly, we consider a function $f(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t)$ defined by:
$f(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t)~=~\mathrm{e}^{K(t)}\theta^{N_{t}}\mathrm{e}^{B(t)\lambda_{t}}~\mathrm{e}^{-\nu C_{t}}\mathrm{e}^{\phi\Lambda_{t}}\psi^{M_{t}},$
@@
Applying Itô’s differentiation rule to $f$ gives:
$f(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t)$ $=$ $f(\lambda_{0},N_{0},C_{0},M_{0},\Lambda_{0},0)+\int^{t}_{0}\delta(a-\lambda_{u})\frac{\partial f}{\partial\lambda}\mathop{}\!\mathrm{d}u+\int^{t}_{0}\lambda_{u}\frac{\partial f}{\partial\Lambda}\mathop{}\!\mathrm{d}u$
$+\int^{t}_{0}\lambda_{u-}\bigg(\int^{\infty}_{0}\int^{\infty}_{0}f(\lambda_{u-}+y,N_{u-}+1,C_{u-}+\xi,M_{u-},\Lambda_{u-},u)$
$\quad\times\mathop{}\!\mathrm{d}G(y)\mathop{}\!\mathrm{d}J(\xi)-f(\lambda_{u-},N_{u-},C_{u-},M_{u-},\Lambda_{u-},u)\bigg)\mathop{}\!\mathrm{d}u$
$+\int^{t}_{0}\rho_{u-}\bigg(\int^{\infty}_{0}f(\lambda_{u-}+x,N_{u-},C_{u-},M_{u-}+1,\Lambda_{u-},u)\mathop{}\!\mathrm{d}H(x)$
$\quad-f(\lambda_{u-},N_{u-},C_{u-},M_{u-},\Lambda_{u-},u)\bigg)\mathop{}\!\mathrm{d}u+{\cal M}_{t},$
where $\{{\cal M}_{t}\}_{t\geq 0}$ is an $(\mathbb{F},\mathbb{P})$-(local)-martingale.
Using Section 2.2.1, we have:
$\hskip-14.22636ptf(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t)$
$=f(\lambda_{0},N_{0},C_{0},M_{0},\Lambda_{0},0)+\int^{t}_{0}{\cal A}f(\lambda_{u-},N_{u-},C_{u-},M_{u-},\Lambda_{u-},u)\mathop{}\!\mathrm{d}u+{\cal M}_{t}.$
If ${\cal A}f(\lambda,n,c,m,\Lambda,t)=0$ then
$f(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t)=f(\lambda_{0},N_{0},C_{0},M_{0},\Lambda_{0},0)+{\cal M}_{t},$
i.e., $\{f(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t)\}_{t\in{\cal T}}$ is an $(\mathbb{F},\mathbb{P})$-(local)-martingale.
Plug $f(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t)$ into the generator in Section 2.2.1 for the standard DCP and set $\mathcal{A}f(\lambda,n,c,m,\Lambda,t)=0$, where
$f(\lambda,n,c,m,\Lambda,t)=\mathrm{e}^{K(t)}\theta^{n}\mathrm{e}^{B(t)\lambda}~\mathrm{e}^{-\nu c}\mathrm{e}^{\phi\Lambda}\psi^{m},$
@@
$0$ $=K^{\prime}(t)+B^{\prime}(t)\lambda+\delta\left(a-\lambda\right)B(t)+\lambda\phi$
@@
$0$ $=\lambda\left\{B^{\prime}(t)-\delta B(t)+\phi+\theta\int_{0}^{\infty}\int_{0}^{\infty}\mathrm{e}^{B(t)y}\mathrm{e}^{-\nu\xi}\mathop{}\!\mathrm{d}G(y)\mathop{}\!\mathrm{d}J(\xi)-1\right\}$
@@
$0$ $=\lambda\bigg\{B^{\prime}(t)-\delta B(t)+\phi+\theta\hat{j}(\nu)\hat{g}(-B(t))-1\bigg\}+K^{\prime}(t)+\delta aB(t)+\rho\left[\psi\hat{h}(-B(t))-1\right],$
@@
    $\mathcal{G}(B):=\int^{B}_{b}\frac{\mathop{}\!\mathrm{d}u}{f_{1}(u)},\qquad B\in\left(b,B^{+}\right),$
@@
$\left.\frac{\mathop{}\!\mathrm{d}f_{1}(B)}{\mathop{}\!\mathrm{d}B}\right|_{B=0}=\ \delta-\theta\hat{j}(\nu)\mu_{G}\ >\ 0.$ (3.5)
@@
$\frac{\mathop{}\!\mathrm{d}^{2}f_{1}(B)}{\mathop{}\!\mathrm{d}B^{2}}=-\theta\hat{j}(\nu)\int^{\infty}_{0}y^{2}\mathrm{e}^{By}\mathop{}\!\mathrm{d}G(y)<0.$
Then $f_{1}(B)$ is strictly concave. This, together with Eqs. 3.4 and 3.5, imply that $f_{1}(B)>0$, for $B\in\left(0,B^{+}\right)$, where $B^{+}>0$ is the smallest positive solution to $f_{1}(B)=0$.
Taking $B=u$ and $t=\tau$, Eqs. 3.2 and 3.3 lead to: